# Constraining accretion-ejection parameters of a newly discovered Galactic transient Swift J151857.0-572147 using IXPE and NuSTAR observations

### v2

@@ 3.3 Joint IXPE and NuSTAR Analysis
| I+N1 | 1447/1325 | ||||||||
| I+N2 | 1213/1144 | ||||||||
| N1 | 991/889 | ||||||||
| N2 | 721/708 | ||||||||
3.4 KERRBB Model
In Fig. 4, we show the model-fitted spectra of Swift J151857 for the joint IXPE and NuSTAR observations. The left and right panels show the spectral fits for the epochs I+N1 and I+N2. The lower panels show the residuals. As we did not have the mass, spin parameters of the BH, and the disk inclination beforehand, we left them free during model fitting, keeping the distance to the source fixed at 10 kpc. However, parameter values may change if the source distance changes significantly. Both data fits returned consistent values of all three parameters in both epochs. The mass () and spin () parameters estimated from the fit vary between and respectively. It possibly indicates that the BH is moderately rotating. The disk inclination () and the mass accretion rate obtained from the fit vary between and ( gm s-1. The low cutoff energy () of keV with a photon index () of indicates a transition to the soft spectral state of the source during the observation epoch after ejecting a high radio flare in the first week of March. The value of the source was high cm-2, which can be due to strong outflows from the outer disk or due to the presence of some blobs along the line of sight that can block the central radiation (see Neilsen & Homan, 2012; Mondal & Jithesh, 2023), requires further study of the outflows from the outer disk. We observed a broad Fe K width which varies between keV. It is expected as the source moved to the soft state, the disk also moved closer to the BH, therefore, relativistic effects can broaden the line (see e.g. Fabian et al., 1989; Iwasawa et al., 1996; Brenneman & Reynolds, 2006; Mondal et al., 2016, and references therein). How close has the disk moved inward closer to the BH? What is the corona geometry (radius and height)? This has been studied in the next section.
@@ 3.5 JeTCAF Model
Fig. 5 shows the JeTCAF model fitted spectra for both epochs. The and values obtained from the fit range between and . The disk moved much closer to the BH with corona radius rg and . However, the higher value of compared to and lower , clearly indicates that the source moved to a soft spectral state. The shock compression ratio () varies between , and falls in the range where the theoretical mass outflow rate is much less (see Chakrabarti, 1999). The mass of the BH () obtained from the fit is and M⊙ for both epochs, which are constant within the error bar and also agrees with KERRBB model fitted BH mass. All model fitted parameters are given in Table 3.
Given the model fitted parameters, we have estimated the mass outflow rate for this source to be (minimum for I+N1 parameters) and (maximum for N2 parameters) using Eq. 16 derived in Chakrabarti (1999). In our model, as discussed earlier, both the corona and the base of the jet behave as a Comptonizing medium. Such a low mass outflow rate also agrees with the soft spectral state.
@@ 4 Conclusions
In this paper, we have performed spectro-polarimetric studies of a recently discovered black hole X-ray binary Swift J151857.0-572147 using simultaneous IXPE and NuSTAR observations during 17 and 18 March 2024. The key findings from this study are summarized below:
The polarization degree and angle of the source are and .
The low cutoff powerlaw energy ( keV) and high powerlaw photon index () indicate a transition to a soft spectral state of the source. This is expected as the source ejected a high radio flare in the first week of March and later moved to the soft spectral state.
Additionally, the high disk mass accretion rate compared to the hot flow rate with lower corona size ( rg) also supports the transition of the source to the soft spectral state.
@@
Our detailed study of the BHXRB source SwiftJ151857 sheds light on the accretion disk properties and the geometry of the corona, as well as the estimation of intrinsic properties of the black hole. Moreover, the joint IXPE and NuSTAR, and independent NuSTAR analysis show consistent results. A detailed timing study is required to further confirm the spectral state of the source.
@@ Acknowledgements
SM and SPS acknowledge Ramanujan Fellowship (# RJF/2020/000113) by SERB-DST, Govt. of India for this research. K.C. acknowledges support from SWIFAR postdoctoral fund of Yunnan University. CBS is supported by the National Natural Science Foundation of China under grant no. 12073021. This research used data products provided by the IXPE Team (MSFC, SSDC, INAF, and INFN) and distributed with additional software tools by the High-Energy Astrophysics Science Archive Research Center (HEASARC), at NASA Goddard Space Flight Center (GSFC). This research has made use of the NuSTAR Data Analysis Software (NuSTARDAS) jointly developed by the ASI Science Data Center (ASDC, Italy) and the California Institute of Technology (Caltech, USA).

### v3

@@ 3.4 KERRBB Model
| I+N1 | 1048/939 | ||||||||
| I+N2 | 890/824 | ||||||||
| N1 | 583/503 | ||||||||
| N2 | 416/388 | ||||||||
The KERRBB model fitted parameters infer that the source is possibly a moderately spinning BH with a low disk inclination. It has an impact on getting low polarization (PD) measurements in IXPE data, as discussed in sec 3.1 and 3.2. If we compare other sources which were observed with high PD in the SS e.g. 4U 1630-47 () and LMC X-3 ( ) are high-inclination, on the contrary, Cyg X-1 and Swift J1727.8-1613 showed low PD having low inclination. Our source falls in the second type and the low PD can be attributed either to the repeated scattering of the trapped radiation in an optically thick medium or the polarization of the disk is perpendicular to the corona in the SS (Krawczynski et al., 2022; Svoboda et al., 2024a, and references therein).
@@ 3.5 JeTCAF Model
Fig. 5 shows the JeTCAF model fitted spectra for both epochs. The and values obtained from the fit range between and . The disk moved much closer to the BH with corona radius rS and . However, the higher value of compared to and lower , clearly indicates that the source moved to a soft spectral state. The shock compression ratio () varies between , and falls in the range where the theoretical mass outflow rate is much less (see Chakrabarti, 1999). The mass of the BH () obtained from the fit ranges - M⊙ for both epochs, which are constant within the error bar and also agrees with KERRBB model fitted BH mass. All model fitted parameters are given in Table 3.
Given the JeTCAF model fitted parameters, we have estimated the mass outflow rate for this source to be (minimum for I+N1 parameters) and (maximum for N2 parameters) using Eq. 16 derived in Chakrabarti (1999). In our model, as discussed earlier, both the corona and the base of the jet behave as a Comptonizing medium. Such a low mass outflow rate also agrees with the soft spectral state.
The JeTCAF model fitted hot accretion flow component () that contributes to the power law emission from the corona is of the cold thermal disk component (). Additionally, the size of the corona () is significantly small (), compared to its size in the intermediate states which is , phenomenologically estimated using timing (QPO) properties (Chatterjee et al., 2024). These estimates infer that the emission is dominated by the thermal disk rather than the corona. Therefore, the geometrical configuration and high disk mass accretion rate may explain that the observed low PD mainly comes from the accretion disk, where the polarization is perpendicular to the disk. Conversely, the PD contribution from the corona is expected to be less due to its smaller size loaded with higher mass, which can trap radiation by repeated scattering resulting in lower PD.
Subsequently, the unpolarized radiation from the central source scattered by the corona of radius and height can gain a net fractional polarization, depending on . Therefore, given the best-fitted corona geometry and , one can estimate the PD using the relation (Begelman & McKee, 1983), , where . The is estimated using Eq. 1 derived in Debnath et al. (2014). For the best-fitted model parameters in epoch I+N2, the above relation predicts expected PD . This range of PD was detected for other sources earlier (see Sec. 1) in their hard or intermediate spectral states, where the corona properties change and that can contribute more to polarization. Therefore, such predictions can be further verified in the future if the source can be observed in other spectral states independently by IXPE and India’s recently launched X-ray Polarimeter Satellite XPoSat, or jointly.
@@ 4 Conclusions
In this paper, we have performed spectro-polarimetric studies of a recently discovered black hole X-ray binary SwiftJ151857 using simultaneous IXPE and NuSTAR observations during 17 and 18 March 2024. The key findings from this study are summarized below:
The first detection of X-ray polarization of the source using IXPE yields PD and PA and in the soft spectral state. The PD value further supports that the changes in spectral states are possibly followed by changes in polarization properties.
The low PD measurement can be due to the low spin and inclination of the disk which agrees with the previous detections. Additionally, in the soft state, the corona has significantly shrunk increasing its optical depth and the emission is coming mainly from the disk. Therefore, it can significantly depolarize the emission by repeated scattering.
The low cutoff powerlaw energy ( keV) and high powerlaw photon index () indicate a transition to a soft spectral state of the source. This is expected as the source ejected a high radio flare in the first week of March and later moved to the soft spectral state. The spectral state has further been confirmed from the non-detection of QPOs using Insight-HXMT observations during this epoch (Chatterjee et al., 2024).
Additionally, the high disk mass accretion rate compared to the hot flow rate with lower corona size ( ) supports the transition of the source to the soft spectral state.
@@
We have detected X-ray polarization of the source SwiftJ151857 for the first time in the soft spectral state and inferred the origin of polarization properties (PD and PA profiles). Our detailed study of the BHXRB source SwiftJ151857 sheds light on the accretion disk properties and the geometry of the corona, as well as the estimation of intrinsic properties of the black hole. The estimated spin parameter and disk inclination can explain the low PD detection. Moreover, the joint IXPE and NuSTAR, and independent NuSTAR analysis show consistent results. A detailed timing study on this source further confirms the spectral state of the source (Chatterjee et al., 2024). We have further predicted the PD using both model-fitted parameters, which is significantly high (4-6%), however, in the range of PD detected for other sources in their hard or intermediate states. Such predictions can be further verified in the future using IXPE and XPoSat, independently or jointly.
@@ Acknowledgements
We thank the referee for making fundamental queries on the physical understanding of the results, which improved the quality of the paper. SM and SPS acknowledge Ramanujan Fellowship (# RJF/2020/000113) by SERB-DST, Govt. of India for this research. K.C. acknowledges support from SWIFAR postdoctoral fund of Yunnan University. CBS is supported by the National Natural Science Foundation of China under grant no. 12073021. This research used data products provided by the IXPE Team (MSFC, SSDC, INAF, and INFN) and distributed with additional software tools by the High-Energy Astrophysics Science Archive Research Center (HEASARC), at NASA Goddard Space Flight Center (GSFC). This research has made use of the NuSTAR Data Analysis Software (NuSTARDAS) jointly developed by the ASI Science Data Center (ASDC, Italy) and the California Institute of Technology (Caltech, USA).